# Compatibility of Kazhdan and Brauer Homomorphism

### v2

@@ 4. Kazhdan isomorphism
and this induces a vector space isomorphism
${\rm Kaz}_{m}^{F}:\mathcal{H}(G_{F},K_{F})\longrightarrow\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}}),$
where $\lambda\in X_{*}(\textbf{T})^{-}$, $(a_{i},a_{j})\in T_{\lambda}(F)$ and $(a_{i}^{\prime},a_{j}^{\prime})$ is the corresponding element in $T_{\lambda}(F^{\prime})$. For $F$ and $F^{\prime}$, sufficiently close fields, the following result is due to Kazhdan ([Kaz86, Theorem A]).
Proposition 4.1.
@@
Given $m\geq 1$. There exists a positive integer $r\geq m$ such that if $F$ and $F^{\prime}$ are $r$-close, the map ${\rm Kaz}_{m}^{F}$ defined above is an algebra isomorphism.
Kazhdan conjectured that one could take $r=m$ in Proposition 4.1, as later proved by R.Ganapathy in her paper [Gan15].
4.2.
@@
Let $F$ and $F^{\prime}$ be the non-Archimedean $m$-close local fields with residue chracteristic $p$. In [Del84], Deligne proved that there is an isomorphism
@@
where $\overline{F}$ (resp. $\overline{F^{\prime}}$) is the seperable closure of $F$ (resp. $F^{\prime}$), and $\mathcal{I}_{F}^{m}$ (resp. $\mathcal{I}_{F^{\prime}}^{m}$) is the $m$-th higher ramification subgroup of the inertia group $\mathcal{I}_{F}$ (resp. $\mathcal{I}_{F^{\prime}}$). Let $E$ be a finite Galois extension of $F$ of prime degree $l$ with $l\neq p$. Following the description of the map ${\rm Del}_{m}$, as mentioned in [Gan15, Section 2.1], we have a finite Galois extension $E^{\prime}$ of $F^{\prime}$ of same degree $l$. Moreover, the extension $E^{\prime}/F^{\prime}$ can be constructed so that $E$ and $E^{\prime}$ are likewise sufficiently close, as ensured by the next lemma.
@@
Let $F$ and $F^{\prime}$ be the $m$-close non-Archimedean local fields, and let $E$ be the finite Galois extension of $F$ of prime degree $l$. Then there exists a finite Galois extension $E^{\prime}$ of $F^{\prime}$ such that $E$ and $E^{\prime}$ are $em$-close, where $e\in\{1,l\}$.
@@
The proof is divided into two parts. In the first part, we consider the case when $E$ is unramified over $F$. The second part deals with totally ramified case. Note that the extension $E/F$ is tamely ramified as $l\neq p$. We denote by $A_{m}$ and $A_{m}^{\prime}$ the quotient rings $\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m}$ and $\mathfrak{o}_{F^{\prime}}/\mathfrak{p}_{F^{\prime}}^{m}$ respectively. Recall that the isomorphism $\Lambda:A_{m}\xrightarrow{\sim}A_{m}^{\prime}$ takes the class of $\varpi$ to that of $\varpi^{\prime}$.
@@ 4.2.1. Unramified case
Let $f^{\prime}$ be an irreducible polynomial in $\mathfrak{o}_{F^{\prime}}[T]$ with its ${\rm mod}-\mathfrak{p}_{F}^{m}$ reduction $\overline{f}^{\prime}$. Let $E^{\prime}$ be the field extension $F^{\prime}[T]/(f^{\prime})$. Then $E^{\prime}$ is unramified over $F^{\prime}$ and there is a ring isomorphism
@@ 4.2.2. Totally ramified case
Suppose $E$ is totally ramified over $F$. Here the field $E$ is of the form $F[T]/(T^{l}-\varpi)$ and its ring of integers $\mathfrak{o}_{E}$ is given by $\mathfrak{o}_{F}[T]/(T^{l}-\varpi)$. Recall that the class of $\varpi$ corresponds to the class of $\varpi^{\prime}$ via the isomorphism $\Lambda$. Now consider the field extension
@@
Then $E^{\prime}$ is totally ramified over $F^{\prime}$ with the ring of integers $\mathfrak{o}_{E^{\prime}}=\mathfrak{o}_{F^{\prime}}[T]/(T^{l}-\varpi^{\prime})$. Let $\overline{\varpi}$ be the image of $\varpi$ under ${\rm mod}-\mathfrak{p}_{F}^{m}$. Then the natural surjection
@@
factorzies through the ideal $\mathfrak{p}_{F}^{m}[T]/(T^{l}-\varpi)$, and it gives the ring isomorphism
@@
Similarly, there is a ring isomorphism
@@
4.2.3.
For each $e\in\{1,l\}$, we fix a ring isomorphism $\Pi:\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em}\rightarrow\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{em}$ as above, induced by $\Lambda$, so that the diagram
@@
is commutative, where the vertical arrows are the natural inclusion maps. We choose uniformizer $\pi$ (resp. $\pi^{\prime}$) of the field $E$ (resp. $E^{\prime}$) such that the class of $\pi$ is mapped to the class of $\pi^{\prime}$ under the isomorphism $\Pi$. For any $\overline{a}\in\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m}$ (or, $\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em}$), we write $\overline{a}^{\prime}$ for the image $\Lambda(\overline{a})$ (or, $\Pi(\overline{a})$), whenever no confusion arises.
@@ 4.3. Compatibility with Galois action
Here, we continue to work with the set up of the preceeding subsection. We fix a generator $\sigma$ of the group ${\rm Gal}(E/F)$. Then $\sigma$ can be regarded as an automorphism of the group $G_{E}$, via its natural action on $G_{E}$. Consider the compact open subgroup
@@
We observe that $K_{E}$ is $\sigma$-plain, and there is an induced action of $\langle\sigma\rangle$ on the Hecke algebra $\mathcal{H}(G_{E},K_{E})$, given by
@@
for all $x\in G_{E}$ and $f\in\mathcal{H}(G_{E},K_{E})$. Similarly, there is an action of $\langle\sigma^{\prime}\rangle$ on the Hecke algebra $\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})$. We choose a group isomorphism $\iota:{\rm Gal}(E/F)\rightarrow{\rm Gal}(E^{\prime}/F^{\prime})$, sending $\sigma$ to a generator, say $\sigma^{\prime}$, that makes commutative the diagram
where the horizontal maps are the ring isomorphisms induced by $\sigma:E\rightarrow E$ and $\sigma^{\prime}:E^{\prime}\rightarrow E^{\prime}$. To be more precise, for any $a\in\mathfrak{o}_{E}$, we have the relation
where $a^{\prime}\in\mathfrak{o}_{E^{\prime}}$ is the preimage of $\overline{a}^{\prime}$ under ${\rm mod}-\mathfrak{p}_{E^{\prime}}^{em}$. We end this section with the following proposition, which shows the compatibility of Kazhdan isomorphism with these Galois actions.
@@
In view of the Cartan decomposition (4.2), it is enough to show that
@@
By Cartan decomposition (4.2), the double coset $K_{E}\sigma(a_{i})\sigma(\pi_{\mu})\sigma(a_{j}^{-1})K_{E}$ is of the form $K_{E}c_{i}\pi_{\mu}c_{j}^{-1}K_{E}$, where $(c_{i},c_{j})\in T_{\mu}(E)$. Then
@@
Recall that the image of $d_{j}$ under the reduction map $\textbf{G}(\mathfrak{o}_{E})\rightarrow\textbf{G}(\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em})$ is denoted by $\overline{d_{j}}$, and $\overline{d_{j}}^{\prime}$ is the image of $\overline{d_{j}}$ under the isomorphism $\textbf{G}(\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em})\xrightarrow{\sim}\textbf{G}(\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{em})$. Let $d_{j}^{\prime}\in\textbf{G}(\mathfrak{o}_{E^{\prime}})$ be a preimage of $\overline{d_{j}}^{\prime}$ under ${\rm mod}-\mathfrak{p}_{E^{\prime}}^{em}$. Then the commutative diagram
@@
$=K_{E^{\prime}}\sigma^{\prime}(a_{i}^{\prime})\sigma^{\prime}(\pi_{\mu}^{\prime})\sigma^{\prime}(a_{j}^{\prime-1})K_{E^{\prime}}.$
This shows that

### v4

@@ 4.2. Kazhdan isomorphism over positive characteristic
$t_{x}\longmapsto t_{\phi(x)},$
which induces an isomorphism of $k$-vector spaces
${\rm Kaz}_{m}^{F}:\mathcal{H}(G_{F},K_{F})\longrightarrow\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}}).$
Kazhdan’s approach in proving the Hecke algebra isomorphism over $\mathbb{C}$, relies on the fact that $\mathcal{H}_{\mathbb{C}}(G_{F},K_{F})$ is finitely presented–which follows from the work of [Ber84]. So, in order to establish that ${\rm Kaz}_{m}^{F}$ is an algebra map, we need the finiteness of Hecke algebras over $k$–which is made available due to seminal work of Dat–Helm–Kurinczuk–Moss ([DHKM24]).
Lemma 4.2.
For any compact open subgroup $K$ of $G_{F}$, the Hecke algebra $\mathcal{H}(G_{F},K)$ is finitely presented.
Proof.
Note that an algebra $\mathcal{A}$ is finitely presented if $\mathcal{A}$ is finitely generated and Noetherian. Then the above lemma follows from [DHKM24, Theorem 1.1 and Corollary 1.4]. ∎
Remark 4.3.
We now recall some results from [Kaz86]. Let $C$ be a finite subset of $X_{*}(\textbf{T})^{-}$ and let
$G_{C}=\bigcup_{\mu\in C}G_{F}(\mu).$
@@
By [Kaz86, Lemma 1.3], there exists a positive integer $n_{C}\geq m$, such that for all $g\in G_{C}$, we have
$gK_{n_{C}}(F)g^{-1}\subseteq K_{F},$
where $K_{n_{C}}(F)$ is the $n_{C}$-th usual congruence subgroup of $G_{F}$. Moreover, if $F$ and $F^{\prime}$ are $n_{C}$-close, then
${\rm Kaz}_{m}^{F}(f_{1}*f_{2})={\rm Kaz}_{m}^{F}(f_{1})*{\rm Kaz}_{m}^{F}(f_{2}),$ (4.4)
for all $f_{1},f_{2}\in\mathcal{H}(G_{F},K_{F})$ supported on $G_{C}$.
The following lemma, which was essential for [Kaz86, Theorem A], is valid over $k$ as well.
Lemma 4.4.
Let $C$ be a finite subset of $X_{*}(\textbf{T})^{-}$. Then
  1. (1)
    for any $\lambda,\mu\in X_{*}(\textbf{T})^{-}$, we have
    $t_{\lambda(\varpi)}*t_{\mu(\varpi)}=t_{(\lambda+\mu)(\varpi)}$
  2. (2)
    for all $x_{1},x_{2}\in G_{C}$, we have
    $t_{x_{1}\lambda(\varpi)x_{2}}=t_{x_{1}}*t_{\lambda(\varpi)}*t_{x_{2}}.$
Proof.
It follows from [All72, Theorem 2] by extension of scalars. ∎
From these, we obtain Kazhdan isomorphism of mod-$l$ Hecke algebras. The proof exactly follows the same line of arguments of [Kaz86, Theorem A].
Theorem 4.5.
Given an integer $m\geq 1$. There exists a positive integer $r\geq m$ such that if $F$ and $F^{\prime}$ are $r$-close, the map ${\rm Kaz}_{m}^{F}$ is an isomorphism of $k$-algebras.
Proof.
We will show that ${\rm Kaz}_{m}^{F}$ preserves the ring structures. Fix a finite subset $C\subseteq X_{*}(\textbf{T})^{-}$ containing 00, such that $C$ generates $X_{*}(\textbf{T})^{-}$ as a semigroup. Consider the finite set
$X_{0}=\bigcup_{\lambda\in C}X_{\lambda},$
where $X_{\lambda}=\big{\{}K_{F}gK_{F}\mid g\in\textbf{G}(\mathfrak{o}_{F})\lambda(\varpi)\textbf{G}(\mathfrak{o}_{F})\big{\}}$. It follows from Lemma 4.4 that the algebra $\mathcal{H}(G_{F},K_{F})$ is generated by the set $\big{\{}t_{g}:g\in X_{0}\big{\}}$, where $t_{g}$ denotes the characteristic function on the double coset $K_{F}gK_{F}$. Assume that $X_{0}=\{x_{1},\dots,x_{s}\}$. Let $k\langle X_{1},\dots,X_{s}\rangle$ denote the non-commutative polynomial algebra (free algebra) with generators $X_{1},\cdots,X_{s}$. Since the Hecke algebra $\mathcal{H}(G_{F},K_{F})$ is finitely presented, there exists a surjective $k$-algebra homomorphism
$\varphi:k\langle X_{1}\dots,X_{s}\rangle\rightarrow\mathcal{H}(G_{F},K_{F})$
$X_{i}\longmapsto t_{x_{i}}$
such that ${\rm Ker}(\varphi)$ is a two-sided ideal generated by non-commutative polynomials $f_{1},f_{2},\dots,f_{d}\in k\langle X_{1},\dots,X_{s}\rangle$. Let
$M={\rm max}\big{\{}a_{i}:1\leq i\leq d\big{\}},$
where $a_{i}$ is the degree of the polynomial $f_{i}$. Then there exists a finite set $B\subseteq\Phi^{+}$ such that
$X_{B}=\bigcup_{\lambda\in B}X_{\lambda}$
contains a compact set of the form $\big{\{}g_{1}g_{2}\dots g_{M}:g_{i}\in X_{0}\big{\}}$. By Remark 4.3, there is an integer $r=r_{B}\geq m$, such that if $F$ and $F^{\prime}$ are $r$-close, we have
${\rm Kaz}_{m}^{F}(f_{1}*f_{2})={\rm Kaz}_{m}^{F}(f_{1})*{\rm Kaz}_{m}^{F}(f_{2}),$
for all functions $f_{1},f_{2}$ supported on $G_{B}$. Now, for each $1\leq i\leq d$, the above identity gives
$f_{i}\big{(}{\rm Kaz}_{m}^{F}(t_{x_{1}}),\dots,{\rm Kaz}_{m}^{F}(t_{x_{s}})\big{)}=0$
or, equivalently
$f_{i}\big{(}t_{\phi(x_{1})},\dots,t_{\phi(x_{s})}\big{)}=0.$
Consider the homomorphism
$\varphi^{\prime}:k\langle X_{1},\dots,X_{s}\rangle\longrightarrow\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})$
$X_{j}\longmapsto t_{\phi(x_{j})}.$
Note that $\varphi^{\prime}$ factors through ${\rm Ker}(\varphi)$, and hence we get a unique algebra morphism
$\widetilde{\varphi^{\prime}}:\mathcal{H}(G_{F},K_{F})\longrightarrow\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})$
such that $\widetilde{\varphi^{\prime}}(t_{x_{j}})=t_{\phi(x_{j})}$, for each $j$. Since $C$ generates $X_{*}(\textbf{T})^{-}$ as a semigroup, using the identities in Lemma 4.4, we get that
$\widetilde{\varphi^{\prime}}(t_{x})=t_{\phi(x)}={\rm Kaz}_{m}^{F}(t_{x}),$
for all $x\in X_{m}(F)$. This shows that $\widetilde{\varphi^{\prime}}={\rm Kaz}_{m}^{F}$, which is an algebra morphism. Hence, the theorem. ∎
4.3. Galois extension of close local fields
@@
Let $F$ and $F^{\prime}$ be two non-Archimedean $m$-close local fields with residue characteristic $p$. Deligne ([Del84, Section 3.5]) proved that there is an isomorphism
@@
where $\overline{F}$ (resp. $\overline{F^{\prime}}$) is the seperable closure of $F$ (resp. $F^{\prime}$), and $\mathcal{I}_{F}^{m}$ (resp. $\mathcal{I}_{F^{\prime}}^{m}$) is the $m$-th higher ramification subgroup of the inertia group $\mathcal{I}_{F}$ (resp. $\mathcal{I}_{F^{\prime}}$). Let $E$ be a finite Galois extension of $F$ of prime degree $l$ with $l\neq p$. Following the description of the isomorphism ${\rm Del}_{m}$ (see [Del84, Section 1.3] and [Gan15, Section 2.1]), we have a finite Galois extension $E^{\prime}/F^{\prime}$ of same degree $l$. Moreover, the extension $E^{\prime}/F^{\prime}$ can be so constructed that the fields $E$ and $E^{\prime}$ are also sufficiently close.
@@
[…]
@@
The lemma follows from [Del84]. For completeness, we give a sketch of the proof here. The proof is in two parts. In the first part, we consider the situation where $E$ is unramified over $F$. The second part deals with totally ramified case. Note that the extension $E/F$ is tamely ramified as $l\neq p$. We denote by $A_{m}$ and $A_{m}^{\prime}$ the quotient rings $\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m}$ and $\mathfrak{o}_{F^{\prime}}/\mathfrak{p}_{F^{\prime}}^{m}$ respectively. Recall that the isomorphism $\Lambda:A_{m}\xrightarrow{\sim}A_{m}^{\prime}$ takes the class of $\varpi$ to that of $\varpi^{\prime}$.
@@ 4.3.1. Unramified case
Let $f^{\prime}$ be an irreducible polynomial in $\mathfrak{o}_{F^{\prime}}[T]$ with its ${\rm mod}-\mathfrak{p}_{F}^{m}$ reduction $\overline{f}^{\prime}$. Let $E^{\prime}$ be the field extension $F^{\prime}[T]/(f^{\prime})$. Then $E^{\prime}$ is unramified over $F^{\prime}$, and there is a ring isomorphism
@@ 4.3.2. Totally ramified case
Suppose $E$ is totally ramified over $F$. Here the field $E$ is of the form $F[T]/(T^{l}-\varpi)$ and its ring of integers $\mathfrak{o}_{E}$ is given by $\mathfrak{o}_{F}[T]/(T^{l}-\varpi)$. Recall that the class of $\varpi$ corresponds to the class of $\varpi^{\prime}$ via the isomorphism $\Lambda$. Now consider the field extension,
@@
Then $E^{\prime}$ is totally ramified over $F^{\prime}$ with the ring of integers $\mathfrak{o}_{E^{\prime}}=\mathfrak{o}_{F^{\prime}}[T]/(T^{l}-\varpi^{\prime})$. Let $\overline{\varpi}$ be the image of $\varpi$ under ${\rm mod}-\mathfrak{p}_{F}^{m}$. Then, the natural surjection
@@
factorizes through the ideal $\mathfrak{p}_{F}^{m}[T]/(T^{l}-\varpi)$, and it gives the ring isomorphism
@@
Similarly, there is a ring isomorphism,
@@
Remark 4.7.
Let $F$, $F^{\prime}$, $E$ and $E^{\prime}$ be as in Lemma 4.6. We simply write $E$ and $E^{\prime}$ are $em$-close, where we define $e=1$ if $E/F$ is unramified and $e=l$ if $E/F$ is totally ramified. We fix a ring isomorphism $\Pi:\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em}\rightarrow\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{em}$ as above, induced by $\Lambda$, which makes commutative the diagram
@@
Here, the vertical arrows are the natural inclusion maps. We choose uniformizer $\pi$ (resp. $\pi^{\prime}$) of the field $E$ (resp. $E^{\prime}$) such that the class of $\pi$ is mapped to the class of $\pi^{\prime}$ under the isomorphism $\Pi$. For any $\overline{a}\in\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m}$ (or, $\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em}$), we write $\overline{a}^{\prime}$ for the image $\Lambda(\overline{a})$ (or, $\Pi(\overline{a})$), whenever no confusion arises.
Remark 4.8.
Let $\Gamma$ and $\Gamma^{\prime}$ denote the Galois groups ${\rm Gal}(E/F)$ and ${\rm Gal}(E^{\prime}/F^{\prime})$ respectively. We choose generators $\sigma$ and $\sigma^{\prime}$ of $\Gamma$ and $\Gamma^{\prime}$ respectively, which makes commutative the following diagram
We now describe the generators $\sigma$ and $\sigma^{\prime}$ in totally ramified case. For the unramified case, one similarly chooses $\sigma$ and $\sigma^{\prime}$. So, assume that $E/F$ is totally ramified. Then, Lemma 4.6 says that $E^{\prime}/F^{\prime}$ is also totally ramified and $E$ is $lm$-close to $E^{\prime}$. Since $l\neq p$, the extension $E/F$ is tamely ramified, and we have
$E=\dfrac{F[T]}{(T^{l}-\varpi)}$ and $E^{\prime}=\dfrac{F^{\prime}[T]}{(T^{l}-\varpi^{\prime})}$.
From (4.3.2) of Lemma 4.6, we get the following isomorphism
$\Pi:\frac{A_{m}[T]}{(T^{l}-\overline{\varpi})}\longrightarrow\frac{A_{m}^{\prime}[T]}{(T^{l}-\overline{\varpi^{\prime}})},$
induced by the ring isomorphism $\Lambda:A_{m}\rightarrow A_{m}^{\prime}$ and it is given by
$\Pi\Big{(}\sum_{i=0}^{l-1}\overline{x_{i}}T^{i}+(T^{l}-\overline{\varpi})\Big{)}=\sum_{i=0}^{l-1}\Lambda(\overline{x_{i}})T^{i}+(T^{l}-\overline{\varpi^{\prime}}),$
where $\overline{x_{i}}\in A_{m}$, for each $i$. Let $\sigma$ and $\sigma^{\prime}$ be the automorphisms of $E$ and $E^{\prime}$ respectively, that sends the class of $T$ to the class of $T^{2}$, i.e.,
$\sigma(T+(T^{l}-\varpi))=T^{2}+(T^{l}-\varpi)$
and
$\sigma^{\prime}(T+(T^{l}-\varpi^{\prime}))=T^{2}+(T^{l}-\varpi^{\prime}),$
Then $\sigma$ (resp. $\sigma^{\prime}$) is a generator of $\Gamma$ (resp. $\Gamma^{\prime}$). Moreover, the automorphism $\sigma$ induces a ring isomorphism $\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em}\xrightarrow{\sim}\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em}$, and the induced map is again denoted by $\sigma$. Similarly for $\sigma^{\prime}$. Then we have
$\Pi\circ\sigma=\sigma^{\prime}\circ\Pi.$
@@ 4.4. Compatibility with Galois action
Here, we continue to work with the setup of the preceding subsection. We choose the generator $\sigma$ (resp. $\sigma^{\prime}$) of $\Gamma$ (resp. $\Gamma^{\prime}$), as discussed in Remark 4.8. We have the following commutative diagram
where $e=1$ if $E/F$ is unramified and $e=l$ if $E/F$ is totally ramified. To be more precise, for any $a\in\mathfrak{o}_{E}$, we have the relation
where $a^{\prime}\in\mathfrak{o}_{E^{\prime}}$ is the preimage of $\overline{a}^{\prime}$ under ${\rm mod}-\mathfrak{p}_{E^{\prime}}^{em}$. We fix an isomorphism $\iota:\Gamma\xrightarrow{\sim}\Gamma^{\prime}$ that sends the generator $\sigma$ to $\sigma^{\prime}$.
4.5.
Note that $\sigma$ can be regarded as an automorphism of the group $G_{E}$ via its natural action on $G_{E}$. Consider the compact open subgroup,
@@
Since $K_{E}$ is a pro-$p$ subgroup of $G_{E}$, it follows from the arguments of [Fen24, Lemma 6.6] that $K_{E}$ is $\sigma$-plain. There is an induced action of $\langle\sigma\rangle$ on the Hecke algebra $\mathcal{H}(G_{E},K_{E})$, given by
@@
for all $x\in G_{E}$ and $f\in\mathcal{H}(G_{E},K_{E})$. Similarly, there is an action of $\langle\sigma^{\prime}\rangle$ on the Hecke algebra $\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})$. We end this section with the following proposition, which shows the compatibility of Kazhdan isomorphism with these Galois actions.
@@
Because of the Cartan decomposition (4.2), it is enough to show that
@@
By Cartan decomposition (4.2), the double coset space $K_{E}\sigma(a_{i})\sigma(\pi_{\mu})\sigma(a_{j}^{-1})K_{E}$ is of the form $K_{E}c_{i}\pi_{\mu}c_{j}^{-1}K_{E}$, for some $(c_{i},c_{j})\in T_{\mu}(E)$. Then
@@
Recall that the image of $d_{j}$ under the reduction map $\textbf{G}(\mathfrak{o}_{E})\rightarrow\textbf{G}(\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em})$ is denoted by $\overline{d_{j}}$, and $\overline{d_{j}}^{\prime}$ is the image of $\overline{d_{j}}$ under the isomorphism $\textbf{G}(\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em})\xrightarrow{\sim}\textbf{G}(\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{em})$. Let $d_{j}^{\prime}\in\textbf{G}(\mathfrak{o}_{E^{\prime}})$ be a preimage of $\overline{d_{j}}^{\prime}$ under ${\rm mod}-\mathfrak{p}_{E^{\prime}}^{em}$. Then, the commutative diagram
@@
$=K_{E^{\prime}}\sigma^{\prime}(a_{i}^{\prime})\sigma^{\prime}(\pi_{\mu}^{\prime})\sigma^{\prime}(a_{j}^{\prime-1})K_{E^{\prime}},$
which shows that